OpenStudy (anonymous):

Integration sin^6(theta)cos^5(theta) dtheta

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin^6(\theta)\cos^5(\theta) d \theta\]

OpenStudy (anonymous):

If I am not wrong, I think first I should change sin^6(theta) to (sin^2(theta))^3 ?

ganeshie8 (ganeshie8):

try changing cos^5 to cos^4 * cos = (1-sin^2)^2 * cos

OpenStudy (anonymous):

so now I should use u = sin correct?

ganeshie8 (ganeshie8):

yes that will do

OpenStudy (anonymous):

I got \[1/11\sin^11(\theta)-2/9\sin^9(\theta)+1/7\sin^7(\theta)+C\]

OpenStudy (anonymous):

1/11 (sin(theta))^11 - 2/9 (sin(theta))^9 +1/7 (sin(theta))^7 +C

ganeshie8 (ganeshie8):

it is simplifying too much

OpenStudy (anonymous):

yeah, after I substitute. I had u^6 (1-u^2)^2

OpenStudy (anonymous):

then u^6 (u^4 -2u^2 +1) after foil

OpenStudy (anonymous):

then u^10 -2u^8 +u^6

ganeshie8 (ganeshie8):

ahh that looks trivial, no need to verify with wolfram ;)

OpenStudy (anonymous):

then antideriving it would be 1/11 u^11 - 2/9 u^9 + 1/7 u^7 +C

OpenStudy (anonymous):

yeah wolfram always does that with trigs :[

ganeshie8 (ganeshie8):

just a dumb tool >.>